From an urn containing 5 red and 5 white balls, an unordered randomsample of size 5 is drawn. Given that there are at least 2 red balls in the sample, find the conditional probability that there are exactly 3 red balls in the sample
Answer by harpazo(655)

Firstly,this is an algebra site. Secondly, tutors like myself provide
free tutoring service not OBLIGATED service. Thirdly, you did not show any work or effort. Fourthly, you are rude.
Lastly, you are lazy.
Here it is lazy:
Ways to get
2 red, 3 white = 5c2*5c3 = 100
3 red, 2 white = 5c3*5c2 = 100
4 red, 1 white = 5c4*5c1 = 25
5 red, 0 whire = 5c5 = 1 ............. total 226
P(exactly 3 red | at least 2 red) = n(3red)/n(at least 2 red)
= 100/226, ≈ 0.4428
Done!
Next time, show some work. How else are we to know your understanding of the topic?